Hey payday loan company charges a $60 fee for a $500 payday loan that will be repaid in 15 days. Treating the fee as interest paid what is the equivalent annual interest rate? round to the nearest percent
1
Expert's answer
2019-05-27T13:03:43-0400

Treating the fee as interest paid and using simple interest formula the equivalent annual interest rate is:

r = I/(P×t) = $60/(500×15/365)×100% = 292% per year.